OpenStudy (anonymous):

I need Help Pleaaase Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (jdoe0001):

hmm take common factor to \(\bf 7sin^2(x)-14sin(x)\) what do you get for the common factor there?

OpenStudy (anonymous):

2 and 7?

OpenStudy (jdoe0001):

or hmm, lemme do it quick so \(\bf 7sin^2(x)-14sin(x)+2=-5 \\ \quad \\\ [7sin^2(x)-14sin(x)]=-7 \\ \quad \\ {\color{brown}{ 7sin(x)}}[sin(x)-2]=-7\) and just solve for sin(x) :)

OpenStudy (anonymous):

What do I do to solve for sin(x)

OpenStudy (jdoe0001):

hmmm lemme hmm actually... lemme redo that some

OpenStudy (anonymous):

Okay

OpenStudy (jdoe0001):

\(\bf 7sin^2(x)-14sin(x)+2=-5 \\ \quad \\\ [7sin^2(x)-14sin(x)]=-7 \\ \quad \\ {\color{brown}{ 7}}[sin^2(x)-2sin(x)]=-7 \\ \quad \\ sin^2(x)-2sin(x)=\cfrac{\cancel{-7}}{\cancel{7}} \\ \quad \\ sin^2(x)-2sin(x)=-1\implies {\color{brown}{ sin}}^2(x)-2{\color{brown}{ sin}}(x)+1=0\) there... notice, it's just a quadratic equation, so, just solve for sin(x), see what you get :)

OpenStudy (anonymous):

Is it just pi?

OpenStudy (jdoe0001):

hmm waht did you get for sin(x) anyway?

OpenStudy (anonymous):

I got pi because I located point (-1, 0) on the unit circle chart. Is that now how you do it?

OpenStudy (anonymous):

not*

OpenStudy (jdoe0001):

hmmm well, I assume you're meant to solve the equation... so , you'd want to solve the quadratic, to see what you get for sin(x) first :)

OpenStudy (anonymous):

Okay so How do I do that? :(

OpenStudy (jdoe0001):

well, is a quadratic notice \(\begin{array}{llll} {\color{brown}{ sin}}^2(x)&-2{\color{brown}{ sin}}(x)&+1&=0\\ {\color{brown}{ x}}^2&-2{\color{brown}{ x}}&+1&=0 \end{array}\) so.. you'd find "the variable" per se, like you'd, on any quadratic it just so happens that in this case, the variable is sin(x)

OpenStudy (jdoe0001):

for example, if you had say... \(\bf x^2-2x+1=0\) what would you factor that to?

OpenStudy (anonymous):

(x-1)^2 + 2(x+1) = 0 ? Or am I way off

OpenStudy (jdoe0001):

hmm well, you're supposed to solve for "x" :)

OpenStudy (jdoe0001):

wouldn't it give you a couple of binomials?

OpenStudy (anonymous):

oh yeah so (x-1)^2=0 and 2(x+1) = 0